# Talk2PC: Enhancing 3D Visual Grounding through LiDAR and Radar Point Clouds Fusion for Autonomous Driving

### v1

@@ Abstract
Embodied outdoor scene understanding forms the foundation for autonomous agents to perceive, analyze, and react to dynamic driving environments. However, existing 3D understanding is predominantly based on 2D Vision-Language Models (VLMs), collecting and processing limited scene-aware contexts. Instead, compared to the 2D planar visual information, point cloud sensors like LiDAR offer rich depth information and fine-grained 3D representations of objects. Meanwhile, the emerging 4D millimeter-wave (mmWave) radar is capable of detecting the motion trend, velocity, and reflection intensity of each object. Therefore, the integration of these two modalities provides more flexible querying conditions for natural language, enabling more accurate 3D visual grounding. To this end, in this paper, we exploratively propose a novel method called TPCNet, the first outdoor 3D visual grounding model upon the paradigm of prompt-guided point cloud sensor combination, including both LiDAR and radar contexts. To adaptively balance the features of these two sensors required by the prompt, we have designed a multi-fusion paradigm called Two-Stage Heterogeneous Modal Adaptive Fusion. Specifically, this paradigm initially employs Bidirectional Agent Cross-Attention (BACA), which feeds dual-sensor features, characterized by global receptive fields, to the text features for querying. Additionally, we have designed a Dynamic Gated Graph Fusion (DGGF) module to locate the regions of interest identified by the queries. To further enhance accuracy, we innovatively devise an C3D-RECHead, based on the nearest object edge. Our experiments have demonstrated that our TPCNet, along with its individual modules, achieves the state-of-the-art performance on both the Talk2Radar and Talk2Car datasets.
@@
3D visual grounding, vision-language model, 4D radar, LiDAR, multi-sensor fusion, autonomous driving
@@
I Introduction
With the rapid advancement of Autonomous Driving (AD) and Intelligent Transportation Systems (ITS), 3D perception based on multi-sensor fusion has been widely adopted in Autonomous Vehicles (AVs), robotics, and roadside perception systems [1, 2, 3]. As a primary sensor for 3D perception, LiDAR provides accurate positioning and detailed 3D representations of objects, and has been proved suitable for numerous perception tasks in AD [4, 5, 6, 7, 8, 9]. However, LiDAR cannot perceive crucial information such as the motion trend and velocity, which are essential for surrounding entities. Luckily, the mmWave radar (radar) is capable of sensing the distance and orientation of each object, with a detection range that exceeds that of LiDAR [10]. Moreover, it can capture motion and velocity, and operates reliably under adverse weather conditions [11]. The latest 4D radar offers richer point cloud data when compared with conventional [12], showing significant potential for 3D perception as well [13, 14, 15, 16, 17, 18, 19]. Consequently, a considerable amount of research has focused on the fusion of LiDAR and radar for enhanced 3D perception [20, 21, 22, 23, 24, 25, 26, 27, 28, 29].
@@ I Introduction
Besides, as Vision-Language Models (VLMs) are progressively applied to interactive perception and embodied intelligence, they enable AVs and robots to not only perceive scenes but also understand human intentions and locate corresponding objects [30, 31]. Nonetheless, the current advancements have predominantly focused on the integration with visual modalities [32], while there has been limited interaction with 3D point cloud sensors, particularly in the domain of language-guided multi-sensor fusion for 3D visual grounding in traffic scenarios. For instance, Talk2Car [33] proposes a benchmark of 2D visual grounding on the image plane from the view of a driving car. WaterVG [34] and NanoMVG [35] focus on the 2D visual grounding based on camera-radar fusion. Cheng et al. [36] extended Talk2Car to LiDAR-based 3D visual grounding while Talk2Radar [37] explores the capacity of 4D radars on 3D visual grounding. Based on the aforementioned, we empirically argue that 2D image plane is inadequate for accurately perceiving the precise spatial coordinates of objects, particularly in terms of depth information. Moreover, the object attributes captured by a single sensor are inherently limited. When the content described by the prompt does not correspond to the information perceived by the sensor, it will lead to incorrect model optimization and parameter learning.
To address the aforementioned issues, in this paper, we propose an exploratory approach that leverages natural language guidance for LiDAR and radar fusion in a dual-sensor framework for 3D visual grounding, as illustrated briefly in Fig. 1. Given that the description patterns of object attributes in prompts are not fixed, it necessitates the dynamic weighting for the importance of the LiDAR and radar in relation to the prompt. Unlike the previous methods that employed static fusion strategies based on convolutions for LiDAR and radar [38, 26, 23, 27, 39], we design a novel Bi-directional Agent Cross Attention (BACA). This approach performs multi-scale feature fusion in the encoder stage, adopting LiDAR and radar as the primary information source for the query alternately. Notably, BACA significantly alleviates the computational complexity associated with vanilla cross-attention while enabling efficient dynamic modeling. Furthermore, to mitigate the issue of false positives from point cloud sensors in 3D visual grounding and to correlate multiple point cloud objects mentioned in the prompt, we propose a novel text-to-point-cloud fusion strategy, termed Dynamic Gated Graph Fusion (DGGF). The limitation of previous Gated Graph Fusion (GGF) [37] is that it uses a static graph construction method, all features construct the same graph, and do not consider the similarity of nodes, which makes it impossible to adjust the graph structure according to the image content, limiting the advantages of graph-based models. In contrast, DGGF constructs a linguistically-conditional candidate region modeling process by utilizing a dynamic axial graph neural network based on feature gating. Furthermore, to prevent prediction errors caused by the absence of object center point clouds and to align the depth information as closely as possible with the descriptions in the prompt, we propose a 3D visual grounding prediction head, termed Corner3D-RECHead (C3D-RECHead). This head uses the region with denser point clouds near the corner of the object closest to the ego vehicle as the anchor centroid.
The main contributions of this paper are summarized as below:
@@
  • We propose a novel model, named TPCNet, which performs 3D visual grounding by leveraging textual semantics to guide dual point cloud sensors. TPCNet is capable of adaptively and dynamically aligning and fusing linguistic features with heterogeneous point cloud data. It achieves state-of-the-art performance on both the Talk2Radar and Talk2Car datasets.
@@
  • We propose the Bi-directional Agent Cross Attention (BACA) to fuse the features of LiDAR and radar with sufficient consideration of respective physical characteristics, which enables TPCNet to embed radar representations of object motion alongside LiDAR’s refined 3D geometric understanding of the environment.
@@
  • We propose Dynamic Gated Graph Fusion (DGGF), which constructs a dynamic axial graph for the adaptive fusion of point cloud features in the lingusitic feature space, while filtering regions of interest.
@@
Figure 2: The architecture of our proposed TPCNet. In TPCNet, Bidirectional Agent Cross Attention (BACA), Dynamic Gated Graph Fusion (DGGF) and C3D-RECHead are three modules proposed by us, which are illustrated in detail in Sub-section III-C, III-D and III-E.
@@ II Related Works
Leveraging the high-precision 3D sensing capabilities of LiDAR and the all-weather robustness and long-range sensing strengths of radar, the fusion of LiDAR and radar has been regarded as a complementary solution for 3D object detection. The works in [26, 23, 40, 38] propose the bi-directional radar-LiDAR fusion module, which is based on the convolution-based attention with the product of feature maps to weigh the respective features dynamically. Besides, in [26], the Interaction-Based Multi-Modal Fusion (IMMF) module is proposed for radar-LiDAR fusion based on bi-directional convolutional weighting. Meng et al. [25] introduced cross fusion to generate the pseudo radar feature upon the guidance of LiDAR. Wang et al. [27] proposed a two-stage LiDAR-radar fusion method to obtain the high-quality composite point cloud features. Building on the above, we observe that, first, several approaches adopt LiDAR-guided radar strategies to complement the features of the two modalities. Second, some methods employ bi-directional fusion to enhance missing features from both modalities while highlighting the object regions. However, these methods are primarily static convolution-based, which demonstrates limited generalization to sensor variations across different scenarios. To address this, the works such as [20, 28] propose attention-based fusion methods to enable dynamic cross-attention and information exchange between modalities. Nonetheless, these attention-based approaches typically operate on high-dimensional features, causing a quadratic increase in computational complexity and significant model parameter overhead, making their representational efficiency compromised. Here, we propose the Bi-directional Agent Cross Attention with linear complexity to take LiDAR and radar as the double information sources of 3D geometric context and motion features to obtain the abundant features that can be aligned with textual prompts.
II-B Outdoor Visual Grounding
@@
In 2D domain, Guan et al. [35] developed a lightweight vision-radar fusion model for waterway monitoring in USVs, while their later work [34] established a multi-task benchmark with two-stage fusion of camera, radar, and language features. However, these approaches remain limited to 2D planes and lack comprehensive 3D spatial reasoning capabilities.
The field has recently evolved toward 3D outdoor grounding, driven by autonomous navigation demands. Cheng et al. [36] pioneered LiDAR-based 3D grounding for driving scenarios, whereas Wu et al. [44] achieved multi-camera object tracking through language prompts and Zhan et al. [45] proposed a monocular-based 3D REC baseline model called Mono3DVG-TR. Radar-based solutions have also emerged, with Guan et al. [37] establishing a 4D radar benchmark on the VoD dataset [46]. Despite these advancements, the current research exhibits three limitations: (1) insufficient integration of complementary sensors, (2) limited adaptability to dynamic textual guidance, and (3) under-explored human-robot interaction paradigms for UGVs.
Our work bridges these gaps through two key contributions. First, we introduce a unified 3D grounding framework to support pluggable point cloud sensors (e.g., LiDAR and radar) with modality-adaptive fusion. Second, we develop dynamic attention mechanisms that automatically adjust sensor weighting based on textual semantics, outperforming conventional sequential fusion approaches. The proposed Talk2PC system demonstrates superior performance in both autonomous driving scenarios and interactive robot navigation tasks.
@@ III-A The Overall Pipeline
Fig. 2 presents the overall architecture of our proposed TPCNet. TPCNet receives inputs from two perception sensors: LiDAR and 4D mmWave radar, both represented in the form of point clouds. Each LiDAR point cloud contains 3D spatial location information $(x,y,z)$ and intensity $(ins)$. Similarly, each 4D mmWave radar point cloud includes 3D spatial location information $(x,y,z)$ aligned to the LiDAR coordinate system, along with radar cross-section $(rcs)$ and compensated radial velocity $(v)$. Additionally, TPCNet accepts textual prompts from the agent to guide the localization of objects in the scene using both sensors.
For the LiDAR and radar inputs, each passes through a pillar-based backbone network, generating multi-scale features at three levels: LiDAR-based features $f^{l}_{i,i\in\{1,2,3\}}$ and radar-based features $f^{r}_{i,i\in\{1,2,3\}}$. The natural language instructions provided by the agent are processed through a text encoder to extract textual features.
The multi-scale LiDAR and radar features are then input into the Bidirectional Agent Cross Attention (BACA) module, where BACA is performed to produce three levels of LiDAR-Radar fusion features $lr_{i,i\in\{1,2,3\}}$. These fused features are further integrated with the textual features in the Dynamic Gated Graph Fusion (DGGF) module, yielding language-embedded contextual features $lc_{i,i\in\{1,2,3\}}$ at three scales.
Subsequently, these features are passed through an FPN to enhance multi-scale feature representation, resulting in the refined features $f_{i,i\in\{1,2,3\}}$. Finally, these features are fed into the C3D-RECHead, which localizes the queried object within the point cloud scene based on the textual instruction, outputting the object’s 3D position, size, and orientation.
@@ III-B Backbones of LiDAR, Radar and Textual Instruction
During the feature encoding stage, the LiDAR and radar point clouds are first processed through pillar-based backbones [47], encoding the point clouds into Bird’s-Eye-View (BEV) pillar features $f^{l}$ and $f^{r}$. Each backbone produces features across three stages, where the feature shape for a single stage is given by $\{f^{l},f^{r}\}\in\mathbb{R}^{C\times H\times W}$.
For the textual instructions provided by the agent, we utilize the text encoder in PointCLIP [48] for encoding instead of standalone text encoders (e.g., BERT [49] or RoBERTa [50]), as these lack an understanding of point cloud information. PointCLIP leverages the powerful contrastive learning capabilities of CLIP [51] to embed point clouds and text into a shared representation space, enabling robust cross-modal representation learning. Using PointCLIP, we obtain the textual features $pl\in\mathbb{R}^{C\times L}$.
@@ III-C Bidirectional Agent Cross Attention
Figure 3: Demonstration of the importance of BACA for 3D visual grounding based on LiDAR and radar.
LiDAR provides high-precision three-dimensional spatial information, while radar excels in diverse weather conditions and is highly effective at detecting the motion of objects. By fusing the data from these two sensors, a more comprehensive environmental understanding can be achieved. Moreover, the textual instructions used to describe the object do not necessarily encapsulate all relevant aspects, such as the semantic meaning, velocity, direction, and depth of the object simultaneously. As such, it is crucial for both LiDAR and radar to dynamically extract features to ensure that the correct attributes are aligned with the textual instructions. Furthermore, for a dual-sensor system, maximizing the utilization of contextual features captured by each sensor and exchanging this information serves to mitigate the information asymmetry inherent in one-way fusion. Additionally, bidirectional fusion facilitates mutual verification between the sensors. Based on these considerations, we propose an efficient and rapid bidirectional fusion module: Bidirectional Agent Cross Attention (BACA).
@@
[…]
where $\mathtt{AdaPool}(\cdot)$ represents the adaptive pooling operation. $l$ is the product result of $h$ and $w$, where $h$ and $w$ denote the height and width of the agent feature. The values of $h$ and $w$ are both smalled than $H$ and $W$. In this context, the size of the agent feature $l$ is a hyperparameter, significantly smaller than the original LiDAR and Radar feature dimensions. It achieves a linear computational complexity of $O(LlC)$. The complexity of our proposed BACA is much smaller than vanilla cross attention’s $O(L^{2}C)$ while still preserving the global cross-modal fusion capability.

### v2

# Enhance 3D Visual Grounding through LiDAR and Radar Point Clouds Fusion for Autonomous Driving

@@ Abstract
Embodied outdoor scene understanding forms the foundation for autonomous agents to perceive, analyze, and react to dynamic driving environments. However, existing 3D understanding is predominantly based on 2D Vision-Language Models (VLMs), which collect and process limited scene-aware contexts. In contrast, compared to the 2D planar visual information, point cloud sensors such as LiDAR provide rich depth and fine-grained 3D representations of objects. Even better the emerging 4D millimeter-wave radar detects the motion trend, velocity, and reflection intensity of each object. The integration of these two modalities provides more flexible querying conditions for natural language, thereby supporting more accurate 3D visual grounding. To this end, we propose a novel method called TPCNet, the first outdoor 3D visual grounding model upon the paradigm of prompt-guided point cloud sensor combination, including both LiDAR and radar sensors. To optimally combine the features of these two sensors required by the prompt, we design a multi-fusion paradigm called Two-Stage Heterogeneous Modal Adaptive Fusion. Specifically, this paradigm initially employs Bidirectional Agent Cross-Attention (BACA), which feeds both-sensor features, characterized by global receptive fields, to the text features for querying. Moreover, we design a Dynamic Gated Graph Fusion (DGGF) module to locate the regions of interest identified by the queries. To further enhance accuracy, we devise an C3D-RECHead, based on the nearest object edge to the ego-vehicle. Experimental results demonstrate that our TPCNet, along with its individual modules, achieves the state-of-the-art performance on both the Talk2Radar and Talk2Car datasets. We release the code at https://github.com/GuanRunwei/TPCNet.
@@
3D visual grounding, LiDAR-radar fusion, interaction perception
I Introduction
With the rapid advancement of Autonomous Driving (AD), 3D perception based on multi-sensor fusion has been widely adopted in Autonomous Vehicles (AVs), robotics, and roadside perception systems [1]. As a primary sensor for 3D perception, LiDAR provides accurate positioning [2] and detailed 3D representations of objects [3], and has been demonstrated to be suitable for numerous perception tasks in AD [4]. However, LiDAR cannot capture crucial information such as the motion trend and velocity [5], which are essential for understanding surrounding entities [6]. In contrast, the mmWave radar (radar) is capable of sensing the distance and orientation of each object, with a detection range that exceeds that of LiDAR [7]. Moreover, radar can capture motion and velocity while operating reliably under adverse weather. The latest 4D radar offers richer point cloud data when compared with conventional 2D radar, showing significant potential for 3D perception as well [8]. Consequently, considerable research has focused on the fusion of LiDAR and radar for enhanced 3D perception [9].
Figure 1: Comparison between camera-based 2D visual grounding and LiDAR-radar-based 3D visual grounding. Under adverse conditions, objects referred to by text prompts containing depth (distance) and motion cues can be localized within the contextual scene using point clouds (LiDAR + radar), which cannot be achieved by cameras.
@@ I Introduction
Besides, as Vision-Language Models (VLMs) are progressively applied to interactive perception and embodied intelligence, they enable AVs and robots not only to perceive scenes but also to understand human intentions and locate corresponding objects [10]. Nonetheless, the current advancements have predominantly focused on the integration with visual modalities, while limited attention has been caught with 3D point cloud sensors, particularly in the domain of language-guided multi-sensor fusion for 3D visual grounding in traffic scenarios. For instance, Talk2Car [11] proposes a benchmark of 2D visual grounding on the image plane from the perspective of a driving car. WaterVG [12] focuses on the 2D visual grounding based on camera-radar fusion. Cheng et al. [13] extended Talk2Car to LiDAR-based 3D visual grounding while Talk2Radar [14] explores the capacity of 4D radars on 3D visual grounding. It is therefore evident that natural language querying of objects offers a more flexible and intuitive interactive paradigm for open-world traffic perception and intelligent transportation systems. However, as illustrated in Fig. 1, we observe that with the increasing demands of environmental perception, the 2D image plane is no longer sufficient for queries involving specific object attributes such as depth, motion trends, or perception under severely adverse conditions. For instance, when querying “a vehicle approximately 20 meters ahead, moving toward the ego-vehicle at a speed of 30 km/h”, an image alone fails to capture any of the object characteristics required by such a natural language prompt. Addressing these queries necessitates accurate 3D scene geometry modeling and a deeper understanding of the physical properties of traffic participants.
To address the aforementioned challenges, an exploratory approach is introduced in this paper that leverages language guidance for LiDAR and radar fusion within a dual-sensor framework for 3D visual grounding, as briefly illustrated in Fig. 2. Given the inherent variability in object attribute descriptions within prompts, a dynamic weighting mechanism is required to adaptively adjust the relative contributions of LiDAR and radar information based on the given prompt.
Unlike existing methods, including Adaptive Fuson Module [15], Query-based Interactive Module [9] and InternRAL [16] that employ static fusion strategies using convolution-based approaches for LiDAR and radar integration, we propose a novel Bi-directional Agent Cross Attention (BACA) mechanism. This approach enables multi-scale feature fusion in the encoder stage, alternating between LiDAR and radar as the primary information source for query processing. Notably, BACA substantially reduces computational complexity compared to conventional cross-attention mechanisms while enabling efficient dynamic modeling.
Furthermore, to mitigate false positives arising from point cloud sensors in 3D visual grounding and to effectively correlate multiple point cloud objects referenced within a prompt, we introduce a Dynamic Gated Graph Fusion (DGGF) strategy. A key limitation of the previous Gated Graph Fusion (GGF) approach [14] lies in its reliance on a static graph construction method, where all features contribute to the same graph without considering node similarity, thereby preventing adaptive graph structure adjustments based on image content. This constraint diminishes the advantages of graph-based models. In contrast, DGGF employs a linguistically conditioned candidate region modeling process, utilizing a dynamic axial graph neural network with feature gating to enhance context-awareness and adaptability.
Additionally, to address prediction errors caused by the absence of object center point clouds and to ensure depth information alignment with prompt descriptions, we propose a 3D visual grounding prediction head, termed Corner3D-RECHead (C3D-RECHead). This head identifies the region with the denser point cloud near the corner of the object closest to the ego vehicle, utilizing this as the anchor centroid to enhance localization accuracy.
The main contributions are summarized as below:
@@
  • We propose TPCNet, which performs 3D visual grounding by leveraging textual semantics to guide dual point cloud sensors. TPCNet adaptively and dynamically aligns and fuses linguistic features with heterogeneous point clouds. It achieves state-of-the-art performance on both the Talk2Radar and Talk2Car datasets.
@@
  • We propose the Bi-directional Agent Cross Attention (BACA) to fuse the features of LiDAR and radar with sufficient consideration of respective physical characteristics, which enables TPCNet to embed radar representations of object motion alongside LiDAR’s 3D geometric understanding of the environment.
@@
  • We propose Dynamic Gated Graph Fusion (DGGF), which constructs a dynamic axial graph for the adaptive fusion of point cloud features in the linguistic feature space, while filtering regions of interest.
@@
The remained content is organized as follows: Section II states the related works; Section III illustrates proposed methods; Section IV demonstrates the experiments; Section V concludes the paper and lists the limitation and future works.
Figure 3: The architecture of proposed TPCNet. In TPCNet, Bidirectional Agent Cross Attention (BACA), Dynamic Gated Graph Fusion (DGGF) and C3D-RECHead are three core modules, which are illustrated in detail in Sub-section III-C, III-D and III-E.
@@ II Related Works
Leveraging the high-precision 3D sensing capabilities of LiDAR and the all-weather robustness and long-range sensing advantages of radar, the fusion of these two modalities has been widely recognized as a complementary approach for 3D object detection. Several studies have introduced bi-directional radar-LiDAR fusion modules, utilizing convolution-based attention mechanisms that compute the product of feature maps to dynamically weight the respective modality features. Among them, Wang et al. [17] propose InterRAL, which fuses the features of LiDAR and radar through the softMax gating mechanism. Yang et al. [9] introduce a query-based interactive feature fusion module to concatenate feature maps of selected LiDAR and radar points. Huang et al. [18] propose Multi-Scale Gated Fusion module to counteract the varying degrees of sensor degradation. Xu et al. [15] propose adaptive fusion module to select salient features of LiDAR and radar. Further advancements in this domain include the cross-fusion approach introduced by Meng et al. [2], which generates pseudo-radar features guided by LiDAR information. Similarly, Wang et al. [16] developed a two-stage LiDAR-radar fusion framework to extract high-quality composite point cloud features.
Building upon these developments, we identify two key trends: (i) several methods adopt LiDAR-guided radar fusion to explore the complementary features of both modalities and (ii) some approaches utilize bi-directional fusion strategies to compensate for missing modality-specific information while simultaneously refining object representations. However, these existing methods predominantly rely on static convolution-based techniques, which exhibit limited generalization capability when applied across varying sensor configurations and environmental conditions.
To address this limitation, recent studies have explored attention-based fusion mechanisms to enable dynamic cross-attention and adaptive information exchange between modalities. Qian et al. [19] propose a dual-branch attention module to dynamically weigh the significance of LiDAR and radar features. Chae et al. [20] introduce the LRF module, which performs LiDAR-based query to match the radar feature by cross attention. Nevertheless, these attention-driven approaches often operate on high-dimensional feature spaces, causing a quadratic increase in computational complexity and a significant model parameter overhead, which compromises their representational efficiency.
To overcome these challenges, we introduce the Bi-directional Agent Cross Attention (BACA) mechanism, which achieves linear computational complexity while leveraging LiDAR and radar as dual sources of 3D geometric context and motion features. This approach ensures the extraction of rich, dynamically aligned features that effectively correspond to textual prompts, enhancing the adaptability and efficiency of multi-modal 3D object detection.
II-B 3D Outdoor Visual Grounding
@@
In 2D domain, [12] established a multi-task benchmark with two-stage fusion of camera, radar, and language features. However, these approaches remain limited to 2D planes and lack comprehensive 3D spatial reasoning capabilities. The field has recently evolved toward 3D outdoor grounding, driven by autonomous navigation demands. Cheng et al. [13] pioneered LiDAR-based 3D grounding for driving scenarios, whereas Zhan et al. [23] proposed a monocular-based 3D REC baseline model called Mono3DVG-TR. Radar-based solutions have also emerged, with Guan et al. [14] establishing a 4D radar benchmark.
Despite these advancements, the current research exhibits three limitations: (1) insufficient integration of complementary sensors, (2) limited adaptability to dynamic textual guidance, and (3) under-explored human-robot interaction paradigms for vehicles. Our work bridges these gaps through two key contributions. First, we introduce a unified 3D grounding framework to support point cloud sensors (e.g., LiDAR and radar) with modality-adaptive fusion. Second, we develop dynamic attention mechanisms that automatically adjust sensor weighting based on textual semantics, outperforming conventional sequential fusion approaches. The proposed Talk2PC system demonstrates superior performance in both autonomous driving scenarios and interactive robot navigation tasks.
@@ III-A The Overall Pipeline
Fig. 3 illustrates the overall architecture of the proposed TPCNet. TPCNet processes inputs from two complementary perception modalities: LiDAR and 4D mmWave radar, both represented as point clouds. Each LiDAR point comprises 3D spatial coordinates $(x,y,z)$ and intensity $(ins)$, while each radar point includes 3D coordinates (aligned to the LiDAR frame), radar cross-section $(rcs)$, and compensated radial velocity $(v)$. In addition, TPCNet receives textual prompts from an agent to guide object localization within the scene.
LiDAR and radar point clouds are independently encoded by a pillar-based backbone, producing multi-scale features at three resolutions: LiDAR features $f^{l}_{i}$ and radar features $f^{r}_{i}$, where $i\in\{1,2,3\}$. Textual instructions are encoded via a language encoder to obtain semantic embeddings.
Multi-scale LiDAR and radar features are fused by the proposed Bidirectional Agent Cross Attention (BACA) module, yielding LiDAR–radar fusion features $lr_{i}$. These are further integrated with textual embeddings through the Dynamic Gated Graph Fusion (DGGF) module to produce language-conditioned contextual features $lc_{i}$.
The resulting features are refined via a Feature Pyramid Network (FPN) to enhance multi-scale representation, producing $f_{i}$. Finally, the proposed C3D-RECHead predicts the queried object’s 3D position, size, and orientation from the fused features, conditioned on the textual prompt.
@@ III-B Backbones of LiDAR, Radar and Textual Instruction
During feature encoding, LiDAR and radar point clouds are independently processed by pillar-based backbones [24], generating Bird’s-Eye-View (BEV) pillar features $f^{l}$ and $f^{r}$. Each backbone produces three-stage feature maps, with each stage represented as $\{f^{l},f^{r}\}\in\mathbb{R}^{C\times H\times W}$. For textual instructions, we adopt the text encoder of PointCLIP [25] rather than standalone language models (e.g., BERT [26]), as the latter lack point cloud awareness. PointCLIP exploits the contrastive learning framework of CLIP [27] to embed both point clouds and text into a unified representation space, facilitating robust cross-modal alignment. This yields textual features $pl\in\mathbb{R}^{C\times L}$.
Figure 4: The significance of BACA for 3D visual grounding based on the fusion of LiDAR and radar.
@@ III-C Bidirectional Agent Cross Attention
As shown in Fig. 4, LiDAR offers high-precision 3D spatial measurements, whereas radar is robust to adverse weather and excels at capturing object motion. Fusing these complementary modalities enables a more complete scene understanding. However, textual descriptions may omit certain attributes, such as semantics, velocity, direction, or depth, making it essential for dual sensors to dynamically extract and align features with the textual instructions. In dual-sensor settings, fully exploiting the context from each modality and enabling reciprocal feature exchange mitigates the asymmetry inherent in unidirectional fusion, while allowing mutual verification between sensors. Motivated by these factors, we introduce the Bidirectional Agent Cross Attention (BACA) module, an efficient mechanism for rapid, bidirectional LiDAR–radar feature fusion conditioned on textual guidance.
@@
$\left\{\begin{aligned} &A_{R}=\mathtt{Flat}(\mathtt{AdaPool}(Q_{R})),A_{R}\in\mathbb{R}^{C\times l},\\ &f_{rc}=\mathtt{Attn}(A_{R},K_{L},V_{L})=\frac{A_{R}K_{L}^{T}}{\sqrt{d}}\cdot V_{L},f_{rc}\in\mathbb{R}^{C\times l},\\ &f_{rm}=\mathtt{Attn}(Q_{R},A_{R},f_{rc})=\frac{Q_{R}A_{R}^{T}}{\sqrt{d}}\cdot f_{rc},f_{rm}\in\mathbb{R}^{C\times L},\\ &f_{rm}=\mathtt{Reshape}(f_{rm}),f_{rm}\in\mathbb{R}^{C\times H\times W},\end{aligned}\right$. (8)
where $\mathtt{AdaPool}(\cdot)$ represents the adaptive pooling operation. $l$ is the product result of $h$ and $w$, where $h$ and $w$ denote the height and width of the agent feature. The values of $h$ and $w$ are both smaller than $H$ and $W$. In this context, the size of the agent feature $l$ is a hyperparameter, significantly smaller than the original LiDAR and Radar feature dimensions. It achieves a linear computational complexity of $O(LlC)$. The complexity of our proposed BACA is much smaller than vanilla cross attention’s $O(L^{2}C)$ while still preserving the global cross-modal fusion capability.